# KNN-MMD: Cross Domain Wireless Sensing via Local Distribution Alignment

### v4

@@ V-G Sensitivity Analysis
Kernels WiGesture [4] WiDar3.0 [48]
@@
Figure 14: One-shot Gesture Recognition: The axis represents the top $p\%$ samples in the testing set used to construct the help set. The dashed and dotted lines indicate the benchmarks for the WiDar3.0 [48] and WiGesture [4] datasets, respectively.
@@ VI Conclusion
[Uncaptioned image] Zijian Zhao received the B.Eng. degree in computer science and technology from School of Computer Since and Engineering, Sun Yat-sen University in 2024. He is currently pursuing the Ph.D. degree in civil engineering (scientific computation) with Department of Civil and Environmental Engineering, The Hong Kong University of Science and Technology. He was a visiting student in Shenzhen Research Institute of Big Data, The Chinese University of Hong Kong (Shenzhen) from 2023 to 2024. His current research interests include deep learning, reinforcement learning, intelligent transportation, mobile computing, and wireless sensing.
[Uncaptioned image] Zhijie Cai received the B.Sc. degree in mathematics and applied mathematics from Sun Yat-sen University in 2022. He is currently pursuing the Ph.D. degree with the Shenzhen Research Institute of Big Data, and the School of Science and Engineering, The Chinese University of Hong Kong, Shenzhen. His current research interests include edge intelligence, wireless sensing, and deep learning.
@@
[Uncaptioned image] Hang Li received the B.E. and M.S. degrees from Beihang University, Beijing, China, in 2008 and 2011, respectively, and the Ph.D. degree from Texas A&M University, College Station, TX, USA, in 2016. He was a postdoctoral research associate with Texas A&M and University of California-Davis (Sept. 2016-Mar. 2018). After being a visiting research scholar (Apr. 2018 – June 2019) at Shenzhen Research Institute of Big Data, Shenzhen, China, he has been a research scientist since June 2019. His current research interests include wireless networks, Internet of things, stochastic optimization, and applications of machine learning. He is recognized as Overseas High-Caliber Personnel (Level C) at Shenzhen in 2020.
[Uncaptioned image] Xiaoyang Li (Member, IEEE) received the B.Eng. degree from the Southern University of Science and Technology (SUSTech) in 2016 and the Ph.D. degree from The University of Hong Kong in 2020. From 2020 to 2022, he was a Presidential Distinguished Research Fellow at SUSTech. He is currently an Associate Research Fellow with the Shenzhen Research Institute of Big Data, and an Adjunct Assistant Professor with the Chinese University of Hong Kong, Shenzhen. He is a recipient of the Youth Talent Support Program by China Communication Society, Forbes China 30 under 30, Overseas Youth Talent in Guangdong, Overseas High-caliber Personnel in Shenzhen, Outstanding Research Fellow in Shenzhen, the Best Paper Award of IEEE 4th International Symposium on Joint Communications & Sensing, and the Exemplary Reviewer of IEEE Wireless Communications Letters. He served as the Editor of Journal of Information and Intelligence, the Workshop/Session Chairs of IEEE ICASSP/ICC/WCNC/PIMRC/MIIS. His research interests include integrated sensing-communication-computation, edge learning, and network optimization.
[Uncaptioned image] Qimei Chen (Member, IEEE) received the Ph.D. degree from the College of Information Science and Electronic Engineering, Zhejiang University, Hangzhou, China, in 2017. She was a Visiting Student with the Department of Electrical and Computer Engineering, University of California at Davis, Davis, CA, USA, from 2015 to 2016. From 2017 to 2022, she was an Associate Researcher with the School of Electric Information, Wuhan University, Wuhan, China, where she has been an Associate Professor, since 2023. Her research interests include intelligent edge communication, unlicensed spectrum, massive MIMO-NOMA, and machine learning in wireless communications. She received the Exemplary Reviewer Certificate of the IEEE Wireless Communications Letters in 2020 and 2023. She has served as a workshop co-chair and TPC Member for IEEE conferences, such as ICC, GLOBECOM, PIMRC, and WCNC. She also served as a Guest Editor for the Special Issues on Heterogeneous Networks of Sensors and NOMA in ISAC (MDPI).

### v5

@@ V-G Sensitivity Analysis
Kernels WiGesture [CSI-BERT] WiDar3.0 [zhang2021widar3]
@@
Figure 14: One-shot Gesture Recognition: The axis represents the top $p\%$ samples in the testing set used to construct the help set. The dashed and dotted lines indicate the benchmarks for the WiDar3.0 [zhang2021widar3] and WiGesture [CSI-BERT] datasets, respectively.
@@ VI Conclusion
[Uncaptioned image] Zijian Zhao (Graduate Student Member, IEEE) received the B.Eng. degree in computer science and technology from School of Computer Since and Engineering, Sun Yat-sen University in 2024. He is currently pursuing the Ph.D. degree in civil engineering (scientific computation) with Department of Civil and Environmental Engineering, The Hong Kong University of Science and Technology. He was a visiting student in Shenzhen Research Institute of Big Data, The Chinese University of Hong Kong (Shenzhen) from 2023 to 2024. His current research interests include deep learning, reinforcement learning, intelligent transportation, mobile computing, and wireless sensing.
[Uncaptioned image] Zhijie Cai (Graduate Student Member, IEEE) received the B.Sc. degree in mathematics and applied mathematics from Sun Yat-sen University in 2022. He is currently pursuing the Ph.D. degree with the Shenzhen Research Institute of Big Data, and the School of Science and Engineering, The Chinese University of Hong Kong, Shenzhen. His current research interests include edge intelligence, wireless sensing, and deep learning.
@@
[Uncaptioned image] Xiaoyang Li (Member, IEEE) is currently an Assistant Professor with Southern University of Science and Technology. He received the Ph.D. degree from The University of Hong Kong. His research interests include integrated sensing-communication-computation and edge learning. He is a recipient of Young Elite Scientists Sponsorship Program by CAST, Forbes China 30 under 30, Young Elite of G20, Overseas Youth Talent in Guangdong, Overseas High-caliber Personnel in Shenzhen, Outstanding Research Fellow in Shenzhen, the Best Paper Award of IEEE 4th International Symposium on Joint Communications and Sensing, the Exemplary Reviewers of IEEE Wireless Communications Letters and Journal of Information and Intelligence (JII). He has served as the Editor of JII, and the Workshop Chairs of IEEE ICASSP/WCNC/PIMRC/MIIS/MediCom.
[Uncaptioned image] Hang Li received the B.E. and M.S. degrees from Beihang University, Beijing, China, in 2008 and 2011, respectively, and the Ph.D. degree from Texas A&M University, College Station, TX, USA, in 2016. He was a postdoctoral research associate with both Texas A&M University and University of California-Davis (Sept. 2016-Mar. 2018). He was a visiting research scholar (Apr. 2018 – June 2019) and a research scientist (June 2019 – May 2025) at Shenzhen Research Institute of Big Data, Shenzhen, China. His current research interests include wireless networks, Internet of things, stochastic optimization, and applications of machine learning. He is recognized as Overseas High-Caliber Personnel (Level C) at Shenzhen in 2020.
[Uncaptioned image] Qimei Chen (Member, IEEE) received the Ph.D. degree from the College of Information Science and Electronic Engineering, Zhejiang University, Hangzhou, China, in 2017. She was a Visiting Student with the Department of Electrical and Computer Engineering, University of California at Davis, Davis, CA, USA, from 2015 to 2016. From 2017 to 2022, she was an Associate Researcher with the School of Electric Information, Wuhan University, Wuhan, China, where she has been an Associate Professor, since 2023. Her research interests include intelligent edge communication, massive MIMO, and machine learning in wireless communications. She received the Exemplary Reviewer Certificate of the IEEE Wireless Communications Letters in 2020 and 2023. She has served as a workshop co-chair and TPC Member for IEEE conferences, such as ICC, GLOBECOM, PIMRC, and WCNC.